OpenStudy (kingogcobraxx):

expand the squared binominal (x+4)^2

OpenStudy (kingogcobraxx):

a. x^2-16 b. x^2+16 c. x^2+8x+16 d. x^2-8x+16

OpenStudy (kingogcobraxx):

c.?

OpenStudy (blank):

yep

OpenStudy (kingogcobraxx):

state the degree of the polynomial xy+3x^2 -7+x

OpenStudy (kingogcobraxx):

a.1 b.2 c.3 d.4

OpenStudy (blank):

The degree of any polynomial is the "x" with the highest exponent which in this case is 3x^2 so the degree is 2.

OpenStudy (kingogcobraxx):

simplify - (a^2-4a-6)-(-a^2-7a-6)

OpenStudy (kingogcobraxx):

2a^2+3a 2a^2+3a-12 2a^2-11a 2a^2+11a

OpenStudy (blank):

I dont think the answer choices match the problem

OpenStudy (kingogcobraxx):

its one of them

OpenStudy (blank):

well i got 11a+12

OpenStudy (kingogcobraxx):

@AaronAndyson @sweetburger

OpenStudy (kingogcobraxx):

@Jadeishere

OpenStudy (jadeishere):

(a^2-4a-6)-(-a^2-7a-6) a^2 - (-a^2) = 2a^2 -4a - (-7a) = -3a -6 - (-6) = 0 2a ^2 - 3a When subtracting polynomial expressions, only subtract the like terms. So a^2, a and no coefficient, in this case :)

OpenStudy (jadeishere):

OH, oops! -4a - (-7a) = 3a, sorry It would actually be 2a^2 + 3a

OpenStudy (blank):

oh I thought the "-" sign in front of the first polynomial was part of the expression so i multiplied that into the polynomial

OpenStudy (jadeishere):

I almost did that, actually.. xD Great job helping him with the other questions. Next time, really, REALLY try to explain the answers a little more in depth :) It helps people even more! Good luck @KingOGCobraXx and thanks @Blank for helping out!

OpenStudy (blank):

thank you for the tip!

OpenStudy (kingogcobraxx):

@Blank ? can you answer my question

OpenStudy (blank):

Since this question asks you to add two polynomials, there is no distributing the minus sign so its a lot easier than a subtraction problem. All you have to do is add the like terms in the two polynomials, so \[2x^2 + 3x^2 = 5x^2\], and \[-11x +11x = 0\], so you get 5x^2-4

OpenStudy (blank):

(3x^2 - 6x + 7) - (x^2 - 6x + 4) 3x^2 - 6x +7 - x^2 + 6x - 4 3x^2 - x^2 - 6x + 6x + 7 - 4 2x^2 + 3 It's the same as before, distribute the minus sign and add/subtract like terms. I'm sure you understand how to do it by now. You should do some on your own for practice now

OpenStudy (kingogcobraxx):

i could do it myself but it would take awhile i have alot of work today and iv taken hours on this test

OpenStudy (blank):

\[-3x^2y^2 * x^4\] Since your multiplying it by x^4 multiply only the x terms so x^2 and x^4. When you are multiplying numbers or variables with exponents you keep the bottom number/variable and add the exponent so you will get x^6. the new expression will be \[-3x^6y^2\]

OpenStudy (blank):

You basically do the same thing. First multiply the x^2 s in the second expression to get \[3x^4y\]. Leave the z as it is because it does not have another like term and multiply all the other numbers/variables. You then get \[18x^5y^4z\]

OpenStudy (blank):

In an expression like that, you multiply the outer exponent outside the parentheses by the exponents inside the parentheses. This one is a simple one, so I would like you to do this one.

OpenStudy (kingogcobraxx):

a or b id say b tho

OpenStudy (blank):

yep its b

OpenStudy (blank):

This one is another simple one. All you have to do is distribute the outside term to each of the inside terms, so that means multiplying 10x to every term inside the parentheses. so it will look like \[(10x)9x^2 + (10x)6x + (10x)7\] So the answer will be \[90x^3 + 60x^2 +70x\]

OpenStudy (blank):

I have to go eat now. Sorry i cant help you any further than this.

OpenStudy (kingogcobraxx):

6 more questions @karim728 ? please

OpenStudy (blank):

sorry i really cant right now

OpenStudy (kingogcobraxx):

@CandyCove

OpenStudy (kingogcobraxx):

@YumYum247